OpenStudy (ksaimouli):

A 60 kg woman stands in an elevator that is accelerating downward at 3.1 m/s^2 Calculate the amount of force she exerts on the elevator floor.

OpenStudy (ksaimouli):

@CliffSedge

OpenStudy (anonymous):

You can subtract 3.1m/s^2 from g and compute the weight.

OpenStudy (ksaimouli):

i am getting 186

OpenStudy (anonymous):

N = m (g-a)

OpenStudy (anonymous):

where a is elevator accelertions and g is gravitation acceleration.,

OpenStudy (anonymous):

You need the total acceleration including g=9.81. Normally the weight would be mg, but the elevator floor is accelerating away, so it's acceleration is taking away from the weight force.

OpenStudy (anonymous):

in these conditions our body feels light

OpenStudy (anonymous):

N = 60 ( 9.8 - 3.1)

OpenStudy (ksaimouli):

i got it (60)(9.8)-(60)(3.1)

OpenStudy (ksaimouli):

402N

OpenStudy (anonymous):

Right. Or factor out the 60 for apparent weight = 60(9.8-3.1). There are only two significant figures, so it's better to round that off to an even 400 N.